I have to find the value of each angle, they are two lines crossing and it gives one angle but not the other, it is in an X shape
Answer by CharlesG2(834)

I have to find the value of each angle, they are two lines crossing and it gives one angle but not the other, it is in an X shape
you will have 4 angles
you know angle A
if B is opposite to A then A and B equal
if C is the supplementary angle to A, and D is the supplementary angle to B, then C = D = 180 - A = 180 - B